OpenStudy (anonymous):

Investigate whether or not it is possible to find numbers s and t which satisfy the possible vector equation. Also, find s and t if it is possible, which in this case it is, but I'm getting wrong values/incorrect answers. (b) \[\left(\begin{matrix}-1 \\ -2\\ 3 \end{matrix}\right)+s\left(\begin{matrix}1 \\ 2\\ -1 \end{matrix}\right)+t\left(\begin{matrix}3 \\ -1\\ 1 \end{matrix}\right)=\left(\begin{matrix}5 \\ 3\\ 4 \end{matrix}\right)\]

OpenStudy (anonymous):

-1+s+3t=5 -2+2s-t=3 3-s+t=4 s+3t =6 ---(1) 2s-t=5 ---(2) s-t=-1 ---(3) (2)-(3): s=6 put this into (3), 6-t=-1 t=7 put t=7 and s=6 into (1) 6+3(7)=6 27=6 which is contradiction Therefore, it is impossible to find s and t

OpenStudy (hoblos):

-1+s +3t = 5 ------(1) -2+2s-t = 3 -------(2) 3-s+t = 4 --------(3) adding (2) and (3) you get 1+s = 7 s=6 then 3-6+t=4 t=7 replacing s and t in (1) -1+6+21 =5 !! false so it is not possible to find s and t

OpenStudy (anonymous):

Hmm... The answer in the text book says 3,1 ....? But i got values 6 and 7 aswell..?

OpenStudy (callisto):

I hope you understand my workings :'(

OpenStudy (anonymous):

Yes, I do. Thanks Callisto... That's the right answer.

OpenStudy (anonymous):

coco is right

OpenStudy (callisto):

I was shocked when they said it's not possible to find s and t...

OpenStudy (callisto):

Wait.. it seems they have a point...

OpenStudy (hoblos):

you get 3and 1 if you solve using equations (1)&(2) but if you use equations (2)&(3) you will get 6 and 7 well in both cases if you replace in the remaining equation it will give you a wrong answer so still the answer is "it is impossible to find s and t"

OpenStudy (anonymous):

Hmm. I got 6 and 7 and 3 and 1, but the textbook says 3 and 1...? So what do I do?

OpenStudy (callisto):

1. write "it is impossible to find s and t" 2. ask your teacher and vice versa

OpenStudy (anonymous):

Hmm, don't have a teacher... just the textbook :P so, I'll just note it as impossible.